# Splat the Net: Radiance Fields with Splattable Neural Primitives

### v1

@@ 1 Introduction
Radiance fields have become a predominant representation for modeling 3D scene appearance. Unlike surface-based approaches, their volumetric formulation is compatible with the gradient-based optimization routines employed during training from multi-view images. Neural representations in particular (Mildenhall et al., 2021) offer unprecedented expressivity in encoding radiance fields. However, rendering an image from a volumetric scene representation is generally challenging: Volume rendering (Kajiya & Von Herzen, 1984) requires the computation of costly integrals along view rays, typically solved using quadrature methods such as ray marching (Max, 1995). As a remedy, primitive-based representations have emerged as an efficient alternative. Popularized by 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), these approaches model radiance fields using a mixture of simple volumetric functions. The key to high rendering efficiency lies in the observation that these primitives can be easily projected onto the image plane, where they become 2D kernels that can be efficiently splatted. A prime example is the 3D Gaussian primitive used in 3DGS, which reduces to a 2D Gaussian splatting kernel (Zwicker et al., 2001). Recently, a variety of functions have been explored as primitives (Mai et al., 2024; von Lützow & Nießner, 2025; Hamdi et al., 2024; Held et al., 2025a; Huang et al., 2024), all relying on relatively simple analytical formulations, which are widely considered essential for enabling efficient conversion into view-dependent splatting kernels.
@@ 2.1 Radiance Field Representation and Rendering
Radiance fields represent the appearance of a 3D scene via a function $F_{\theta}:\left(\mathbf{x},\mathbf{d}\right)\rightarrow\left(\sigma,\mathbf{c}\right)$, which maps a spatial location $\mathbf{x}\in\mathbb{R}^{3}$ and view direction $\mathbf{d}\in\mathbb{S}^{2}$ to a volumetric density $\sigma\in\mathbb{R}_{+}$ and an RGB color $\mathbf{c}\in\mathbb{R}^{3}$. Synthesizing an image from a radiance field involves emission–absorption volume rendering (Kajiya & Von Herzen, 1984) along view rays $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d},$ where $\mathbf{o}\in\mathbb{R}^{3}$ denotes the camera center, and $t\in\mathbb{R}_{+}$ parameterizes the ray. Specifically, each RGB pixel color $C\in\mathbb{R}^{3}$ is computed by evaluating the radiance field along its corresponding view ray:
@@
Neural Radiance Fields (NeRFs) (Mildenhall et al., 2021) represent $F_{\theta}$ using a neural network. Numerous follow-up works extended and enhanced its capabilities (e.g.,  (Barron et al., 2021; 2022; Martin-Brualla et al., 2021)). However, rendering an image from it involves ray marching, that is, discretizing Eq. 1 into a finite sum of samples along each ray (Max, 1995) (red dots in Fig. 2b):
@@
The 3D Gaussian is not the only primitive allowing efficient (approximate) conversion into a 2D splatting kernel $\alpha$ . Recent work has explored a variety of primitive shapes (Mai et al., 2024; von Lützow & Nießner, 2025; Hamdi et al., 2024; Held et al., 2025a; Chen et al., 2024; Li et al., 2024; Gu et al., 2024; Hamdi et al., 2024; Talegaonkar et al., 2025; Liu et al., 2025), all relying on hand-crafted analytic kernels for efficient evaluation of Eq. 3. In contrast, our representation introduces splattable neural primitives, greatly increasing modeling flexibility. Neural components have also been used in the context of primitive splatting, for example, by injecting structure into the representation (Lu et al., 2024) or enforcing spatial regularization (Mihajlovic et al., 2024), yet these methods ultimately splat Gaussian functions. To the best of our knowledge, we are the first to represent the volumetric kernel itself – i.e., the density distribution – as a neural network, making the primitive fundamentally neural rather than merely Gaussian with neural augmentations.
@@ 3.1 Representation
We parameterize $F_{\theta}$ as a mixture of volumetric primitives $\left\{P_{i}\right\}$. Each primitive occupies a volume bounded by an ellipsoid (Mai et al., 2024), which we denote as $B$ (Fig. 3a). This ellipsoid is defined by a center $\mathbf{x}_{B}\in\mathbb{R}^{3}$, a scaling vector $\mathbf{s}_{B}\in\mathbb{R}^{3}$ along its principal axes, and a rotation quaternion $\mathbf{q}_{B}\in\mathbb{R}^{4}$. In accordance with the radiance field formalism, each primitive must define a spatially varying density $\sigma$ and a view-dependent color $\mathbf{c}$ , described next. We define a density field $\sigma(\mathbf{x}):B\rightarrow\mathbb{R}$ within the volume of the ellipsoid as
@@
Here, $W_{1}\in\mathbb{R}^{N_{\sigma}\times 3}$ and $W_{2}\in\mathbb{R}^{1\times N_{\sigma}}$ are weight matrices, while $\mathbf{b}_{1}\in\mathbb{R}^{N_{\sigma}}$ and $\mathbf{b}_{2}\in\mathbb{R}$ are biases. Similar to (Sitzmann et al., 2020), we use a fixed boosting frequency $\omega_{0}$ , which yields a stable initialization. The network structure of Eq. 6 admits an interpretation analogous to a Fourier series, where $W_{1}$ and $\mathbf{b}_{1}$ correspond to frequencies and phases, and $W_{2}$ and $\mathbf{b}_{2}$ are amplitudes and offsets. The normalization by $\mathbf{x}_{B}$ and $\mathbf{s}_{B}$ in Eq. 5 ensures that $f_{\sigma}$ operates on a centered and uniformly scaled domain. In the appendix, we show proof-of-concept extensions of this model to higher-dimensional inputs, including time, which can be easily incorporated into our model by augmenting the network’s input dimensions. To represent view-dependent color, we adopt the Spherical Harmonics basis.
@@ Discussion
We emphasize the efficiency of evaluating the splatting kernel via Eq. 9, which computes a closed-form integral along arbitrary view rays through the neural density field, thereby avoiding the computational cost of ray marching. In contrast to splatting-based Gaussian rendering, relying on an affine approximation of the projection operator (Heckbert, 1989; Zwicker et al., 2004), our method yields perspectively accurate results. Note that the density $\sigma$ in Eq. 5 is never evaluated directly, neither during training nor during view synthesis. Instead, all computations operate directly on its antiderivative $S$ . Yet, in contrast to a light-field-style approach that directly regresses integrated appearance (Sitzmann et al., 2021), our method achieves multi-view consistency by construction.
@@ 3.3 Implementation Details
We initialize $W_{1}\sim\mathcal{U}\left(-1/3,1/3\right)$ and $W_{2}\sim\mathcal{U}\left(-\sqrt{6/N_{\sigma}}/\omega_{0},\sqrt{6/N_{\sigma}}/\omega_{0}\right)$ following (Sitzmann et al., 2020). We set the number of hidden neurons $N_{\sigma}$ to $8$ and the frequency multiplier $\omega_{0}$ to $30$. Similar to 3DGS, we employ four bands of Spherical Harmonics coefficients for color representation. Each neural primitive in our system consists of $99$ parameters in total, around $1.6\times$ more than Gaussian primitives used in 3DGS. We provide a detailed analysis of network configurations in Sec. 5.
@@
We follow the same loss function as 3DGS, and introduce a geometric regularization term to penalize the extreme anisotropy in primitive shapes by minimizing the standard deviation of the components of the scale vector $\mathbf{s}_{B}$ . The effectiveness of this regularization is demonstrated in  Sec. 5. We implement all frameworks in PyTorch (Paszke et al., 2017) and CUDA. All models are trained on a single NVIDIA A40 GPU and evaluated on an NVIDIA RTX 4090 for performance analysis. Due to the complex optimization landscape of neural fields, the convergence of our representation is slower than a Gaussian-based one. We therefore extend training to $100k$ iterations. Additional training details are provided in Appendix A.
@@ 4 Evaluation
In this section, we perform a comprehensive evaluation of neural primitives on novel-view synthesis tasks. We first demonstrate the expressivity of the neural density field (Sec. 4.1). We then perform quantitative and qualitative analysis on synthetic datasets (Sec. 4.2) and real datasets (Sec. 4.3). Please refer to Appendix B for additional results.
@@ 4.1 Primitive Expressivity
Leveraging a flexible neural density field and analytically exact integration, our method faithfully reproduces complex geometries with a small number of primitives. To demonstrate this, we optimize varying numbers of neural and Gaussian primitives to approximate the density fields of several 3D geometries from multiple views. We visualize both the renderings and the color-coded primitives in Fig. 4. We observe that a few neural primitives suffice to represent complex and diverse geometries, such as the teapot’s curved handle, the smooth cut in the leaf, and the triangular leaf petiole. In contrast, Gaussian primitives are limited by their symmetric ellipsoidal shape and soft boundaries, making them unsuitable for accurately representing complex solid structures. Neural primitives achieve superior performance while using $4\times$ fewer primitives and $16\times$ fewer parameters than Gaussian primitives.
@@ 4.2 Synthetic Scenes
We compare our method with 3DGS on the Synthetic NeRF dataset (Mildenhall et al., 2021) across varying memory budgets. Specifically, we resample the original meshes to target vertex counts ($200$, $500$, $1\mathrm{k}$, $2\mathrm{k}$, $5\mathrm{k}$, $10\mathrm{k}$, $20\mathrm{k}$) and use them to initialize primitive positions for optimization, omitting primitive densification. We also include an “unlimited” setting, in which training follows the standard densification procedure with no primitive budget.
@@ Results
We report numerical results in Tab. 1. Our method outperforms 3DGS under limited memory budgets and achieves performance comparable to 3DGS when no memory constraints are imposed. Visual comparisons are shown in Fig. 5. For Ficus, a single primitive can already reconstruct an entire leaf (highlighted by the white frame). Similarly, in Lego, neural primitives capture diverse geometries, such as the front shovel and the rear wheel. In contrast, Gaussian primitives perform poorly on these complex structures, particularly under tight budgets.
@@ 4.3 Real Scenes
For evaluation on real scenes, we follow established practice and use two scenes from Deep Blending (Hedman et al., 2018), two from Tanks & Temples (Knapitsch et al., 2017), and all scenes from the Mip-NeRF360 dataset (Barron et al., 2022). We compare against three method families: (i) splatting-based approaches with analytic primitives – 3DGS (Kerbl et al., 2023), GES (Hamdi et al., 2024), ConvSplat (Held et al., 2025b), BetaGS (Liu et al., 2025), and Vol3DGS (Talegaonkar et al., 2025); (ii) T-3DGS (Mallick et al., 2024), which provides a more sophisticated mechanism for controlling the memory footprint of primitive-based representations; and (iii) monolithic representations – Plenoxels (Fridovich-Keil et al., 2022), INGP (Müller et al., 2022), and MipNeRF360 (Barron et al., 2022). All experiments use the official code released by the respective authors. Since our reproduced baseline results closely match those reported in the respective papers, we report the original numbers for consistency. For a fair comparison, all inference FPS values are measured on a single NVIDIA GeForce RTX 4090 GPU.
@@ 5.1 Neural Integration
Figure 8: We visualize the effect of network width $N_{\sigma}$ and frequency multiplier $\omega_{0}$ on the expressivity of our neural density field.
5.2 Network Width
We analyze how the number of hidden neurons ($N_{\sigma}$ ) and the frequency multiplier ($\omega_{0}$ ) affect the expressivity of our neural representation. We vary $N_{\sigma}\in{4,8,16,32}$ and $\omega_{0}\in{1,10,30}$, and run experiments on Snowflake using a single primitive and Leaf using eight primitives. As shown in Fig. 8, larger $N_{\sigma}$ and higher $\omega_{0}$ better reproduce the Snowflake structure and the smooth contours of Leaf. We further evaluate on real scenes from MipNeRF360, disabling densification and optimizing the same number of primitives with varying $N_{\sigma}$ . With $N_{\sigma}=4$, we obtain an average PSNR of $26.96$; $N_{\sigma}=8$ and $N_{\sigma}=16$ yield $27.21$ and $27.29$, respectively. Although $N_{\sigma}=16$ offers greater expressivity than $8$ in toy settings, this advantage diminishes on real scenes due to the difficulty of optimizing a highly under-constrained problem. Balancing memory footprint and expressivity, we set $N_{\sigma}=8$ and $\omega_{0}=30$ as the default configuration for all experiments.

### v2

@@ 1 Introduction
Radiance fields have become a predominant representation for modeling 3D scene appearance. Unlike surface-based approaches, their volumetric formulation is compatible with the gradient-based optimization routines employed during training from multi-view images. Neural representations, particularly Neural Radiance Fields introduced by Mildenhall et al. (2020), offer unprecedented expressivity in encoding radiance fields. However, rendering an image from a volumetric scene representation is generally challenging: Volume rendering (Kajiya and Von Herzen, 1984) requires the computation of costly integrals along view rays, typically solved using quadrature methods such as ray marching (Max, 1995). As a remedy, primitive-based representations have emerged as an efficient alternative. Popularized by 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), these approaches model radiance fields using a mixture of simple volumetric functions. The key to high rendering efficiency lies in the observation that these primitives can be easily projected onto the image plane, where they become 2D kernels that can be efficiently splatted. A prime example is the 3D Gaussian primitive used in 3DGS, which reduces to a 2D Gaussian splatting kernel (Zwicker et al., 2001). Recently, a variety of functions have been explored as primitives (Mai et al., 2025; von Lützow and Nießner, 2025; Hamdi et al., 2024; Held et al., 2025; Huang et al., 2024), all relying on relatively simple analytical formulations, which are widely considered essential for enabling efficient conversion into view-dependent splatting kernels.
@@ 2.1 Radiance Field Representation and Rendering
Radiance fields represent the appearance of a 3D scene via a function $F_{\theta}:\left(\mathbf{x},\mathbf{d}\right)\rightarrow\left(\sigma,\mathbf{c}\right)$, which maps a spatial location $\mathbf{x}\in\mathbb{R}^{3}$ and view direction $\mathbf{d}\in\mathbb{S}^{2}$ to a volumetric density $\sigma\in\mathbb{R}_{+}$ and an RGB color $\mathbf{c}\in\mathbb{R}^{3}$. Synthesizing an image from a radiance field involves emission–absorption volume rendering (Kajiya and Von Herzen, 1984) along view rays $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d},$ where $\mathbf{o}\in\mathbb{R}^{3}$ denotes the camera center, and $t\in\mathbb{R}_{+}$ parameterizes the ray. Specifically, each RGB pixel color $C\in\mathbb{R}^{3}$ is computed by evaluating the radiance field along its corresponding view ray:
@@
Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) represent $F_{\theta}$ using a neural network. Numerous follow-up works  (Barron et al., 2021; 2022; Martin-Brualla et al., 2021) extended and enhanced its capabilities. However, rendering an image from it involves ray marching, that is, discretizing Eq. 1 into a finite sum of samples along each ray (Max, 1995) (red dots in Fig. 2b):
@@
The 3D Gaussian is not the only primitive allowing efficient (approximate) conversion into a 2D splatting kernel $\alpha$ . Recent work has explored a variety of primitive shapes (Mai et al., 2025; von Lützow and Nießner, 2025; Hamdi et al., 2024; Held et al., 2025; Chen et al., 2024; Li et al., 2025; Gu et al., 2024; Hamdi et al., 2024; Talegaonkar et al., 2025; Liu et al., 2025), all relying on hand-crafted analytic kernels for efficient evaluation of Eq. 3. In contrast, our representation introduces splattable neural primitives, greatly increasing modeling flexibility. Neural components have also been used in the context of primitive splatting, for example, by injecting structure into the representation (Lu et al., 2024) or enforcing spatial regularization (Mihajlovic et al., 2024), yet these methods ultimately splat Gaussian functions. To the best of our knowledge, we are the first to represent the volumetric kernel itself – i.e., the density distribution – as a neural network, making the primitive fundamentally neural rather than merely Gaussian with neural augmentations.
@@ 3.1 Representation
We parameterize $F_{\theta}$ as a mixture of volumetric primitives $\left\{P_{i}\right\}$. Each primitive occupies a volume bounded by an ellipsoid (Mai et al., 2025), which we denote as $B$ (Fig. 3a). This ellipsoid is defined by a center $\mathbf{x}_{B}\in\mathbb{R}^{3}$, a scaling vector $\mathbf{s}_{B}\in\mathbb{R}^{3}$ along its principal axes, and a rotation quaternion $\mathbf{q}_{B}\in\mathbb{R}^{4}$. In accordance with the radiance field formalism, each primitive must define a spatially varying density $\sigma$ and a view-dependent color $\mathbf{c}$ , described next. We define a density field $\sigma(\mathbf{x}):B\rightarrow\mathbb{R}$ within the volume of the ellipsoid as
@@
Here, $W_{1}\in\mathbb{R}^{N_{\sigma}\times 3}$ and $W_{2}\in\mathbb{R}^{1\times N_{\sigma}}$ are weight matrices, while $\mathbf{b}_{1}\in\mathbb{R}^{N_{\sigma}}$ and $\mathbf{b}_{2}\in\mathbb{R}$ are biases. Similar to Sitzmann et al. (2020), we use a fixed boosting frequency $\omega_{0}$ , which yields a stable initialization. The network structure of Eq. 6 admits an interpretation analogous to a Fourier series, where $W_{1}$ and $\mathbf{b}_{1}$ correspond to frequencies and phases, and $W_{2}$ and $\mathbf{b}_{2}$ are amplitudes and offsets. The normalization by $\mathbf{x}_{B}$ and $\mathbf{s}_{B}$ in Eq. 5 ensures that $f_{\sigma}$ operates on a centered and uniformly scaled domain. In the appendix, we also show proof-of-concept extensions of this model to higher-dimensional inputs, including time, which can be easily incorporated into our model by augmenting the network’s input dimensions. To represent view-dependent color, we adopt the Spherical Harmonics basis.
@@ Discussion
We emphasize the efficiency of evaluating the splatting kernel via Eq. 9, which computes a closed-form integral along arbitrary view rays through the neural density field, thereby avoiding the computational cost of ray marching. In contrast to splatting-based Gaussian rendering, relying on an affine approximation of the projection operator (Heckbert, 1989; Zwicker et al., 2004), our method yields perspectively accurate results. Note that the density $\sigma$ in Eq. 5 is never evaluated directly, neither during training nor during view synthesis. Instead, all computations operate directly on its antiderivative $S$ . Yet, in contrast to a light-field-style approach that directly regresses integrated appearance (Sitzmann et al., 2021), our method achieves multi-view consistency by construction. A detailed analysis of the computational cost of the integration procedure is provided in Appendix A.
@@ 3.3 Implementation Details
We initialize $W_{1}\sim\mathcal{U}\left(-1/3,1/3\right)$ and $W_{2}\sim\mathcal{U}\left(-\sqrt{6/N_{\sigma}}/\omega_{0},\sqrt{6/N_{\sigma}}/\omega_{0}\right)$ following Sitzmann et al. (2020). We set the number of hidden neurons $N_{\sigma}$ to $8$ and the frequency multiplier $\omega_{0}$ to $30$. Similar to 3DGS, we employ four bands of Spherical Harmonics coefficients for color representation. Each neural primitive in our system consists of $99$ parameters in total, around $1.6\times$ more than Gaussian primitives used in 3DGS. We provide a detailed analysis of network configurations in Sec. 5.
@@
We follow the same loss function as 3DGS, and introduce a geometric regularization term to penalize the extreme anisotropy in primitive shapes by minimizing the standard deviation of the components of the scale vector $\mathbf{s}_{B}$ . The effectiveness of this regularization is demonstrated in  Sec. 5. We implement all frameworks in PyTorch (Paszke et al., 2017) and CUDA. All models are trained on a single NVIDIA A40 GPU and evaluated on an NVIDIA RTX 4090 for performance analysis. Due to the complex optimization landscape of neural fields, the convergence of our representation is slower than a Gaussian-based one. We therefore extend training to $100k$ iterations. Additional training details are provided in Appendix B.
@@ 4 Evaluation
In this section, we perform a comprehensive evaluation of neural primitives on novel-view synthesis tasks. We first demonstrate the expressivity of the neural density field (Sec. 4.1). We then perform quantitative and qualitative analysis on synthetic datasets (Sec. 4.2) and real datasets (Sec. 4.3) using standard evaluation metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM)  (Wang et al., 2004), and Learned Perceptual Image Patch Similarity (LPIPS)  (Zhang et al., 2018). Please refer to Appendix C for additional results.
@@ 4.1 Primitive Expressivity
Leveraging a flexible neural density field and analytically exact integration, our method faithfully reproduces complex geometries with a small number of primitives. To demonstrate this, we optimize varying numbers of neural and Gaussian primitives to approximate the density fields of several 3D geometries from multiple views. We visualize both the renderings and the color-coded primitives in Fig. 4. We observe that a few neural primitives suffice to represent complex and diverse geometries, such as the teapot’s curved handle, the smooth cut in the leaf, and the triangular leaf petiole. In contrast, Gaussian primitives are limited by their symmetric ellipsoidal shape and soft boundaries, making them unsuitable for accurately representing complex solid structures. For the toy examples of teapot and leaf in Fig. 4, neural primitives achieve superior performance while using $16\times$ fewer primitives and $4\times$ fewer parameters than 3DGS.
@@ 4.2 Synthetic Scenes
We compare our method with 3DGS on the Synthetic NeRF dataset (Mildenhall et al., 2020) across varying memory budgets. Specifically, we resample the original meshes to target vertex counts ($200$, $500$, $1\mathrm{k}$, $2\mathrm{k}$, $5\mathrm{k}$, $10\mathrm{k}$, $20\mathrm{k}$) and use them to initialize primitive positions for optimization, omitting primitive densification. We also include an “unlimited” setting, in which training follows the standard densification procedure with no primitive budget.
@@ Results
We report numerical results in Tab. 1. Our method outperforms 3DGS under limited memory budgets and achieves performance comparable to 3DGS when no memory constraints are imposed. Visual comparisons are shown in Fig. 5. For Ficus, a single primitive can already reconstruct an entire leaf (highlighted by the white frame). Similarly, in Lego, neural primitives capture diverse geometries, such as the front shovel and the rear wheel. In contrast, Gaussian primitives perform poorly on these complex structures, particularly under tight budgets. We further analyze the impact of the primitive initialization strategy. Neural primitives with random initialization achieve an average PSNR of 33.36, which is comparable to 3DGS with random initialization (33.32) and to neural primitives with mesh-based initialization (33.40). This observation indicates that our representation behaves similarly to 3DGS across different initialization strategies.
@@ 4.3 Real Scenes
For evaluation on real scenes, we follow established practice and use two scenes from Deep Blending (Hedman et al., 2018), two from Tanks & Temples (Knapitsch et al., 2017), and all scenes from the Mip-NeRF360 dataset (Barron et al., 2022). We compare against three method families: (i) splatting-based approaches with analytic primitives – 3DGS (Kerbl et al., 2023), GES (Hamdi et al., 2024), ConvSplat (Held et al., 2025), BetaGS (Liu et al., 2025), and Vol3DGS (Talegaonkar et al., 2025); (ii) T-3DGS (Mallick et al., 2024), which provides a more sophisticated mechanism for controlling the memory footprint of primitive-based representations; and (iii) monolithic representations – Plenoxels (Fridovich-Keil et al., 2022), INGP (Müller et al., 2022), and MipNeRF360 (Barron et al., 2022). All experiments use the official code released by the respective authors. Since our reproduced baseline results closely match those reported in the respective papers, we report the original numbers for consistency. For a fair comparison, all inference FPS values are measured on a single NVIDIA GeForce RTX 4090 GPU.
@@ 5 Ablation Studies
Figure 8: We visualize the effect of network configuration ($N_{\sigma}$ and $\omega_{0}$ ) on the expressivity of our neural density field.
@@ 5.1 Neural Integration
Table 3: Ablation study of network configuration ($N_{\sigma}$ and $\omega_{0}$ ) on all MipNeRF360 scenes. For each setup, we report the memory footprint (in MB) and image quality (PSNR).
$N_{\sigma}$ Mem PSNR
$\omega_{0}=1$ $\omega_{0}=10$ $\omega_{0}=30$ $\omega_{0}=50$
4 73 26.35 27.29 27.33 27.12
8 92 26.32 27.42 27.54 27.41
16 129 26.19 27.46 27.62 27.55
5.2 Primitive Network Configuration
We analyze how the number of hidden neurons ($N_{\sigma}$ ) and the frequency multiplier ($\omega_{0}$ ) affect the expressivity of our neural representation. We vary $N_{\sigma}\in\{4,8,16,32\}$ and $\omega_{0}\in\{1,10,30\}$, and run experiments on Snowflake using a single primitive and Leaf using eight primitives. As shown in Fig. 8, larger $N_{\sigma}$ and higher $\omega_{0}$ better reproduce the Snowflake structure and the smooth contours of Leaf. Although $N_{\sigma}=16$ offers greater expressivity than $8$ in toy settings, this advantage diminishes on real scenes due to the difficulty of optimizing a highly under-constrained problem.
We further analyze the effect of the network configuration on all MipNeRF360 scenes. To facilitate a meaningful comparison, we resume training from a pretrained checkpoint and disable densification, thereby eliminating the influence of the number of primitives. We report PSNR values and memory footprint in Tab. 3. As $\omega_{0}$ increases from 0 to 50, the PSNR initially improves and reaches its peak for $\omega_{0}=30$, after which it slightly declines. At $\omega_{0}$ = 30, increasing $N_{\sigma}$ from 4 to 8 yields a larger quality improvement than increasing it from 8 to 16. However, the memory footprint from 8 to 16 is significantly higher (37 MB) than that from 4 to 8 (19 MB). Balancing memory footprint and expressivity, we set $N_{\sigma}=8$ and $\omega_{0}=30$ as the default configuration for all experiments.